OpenStudy (anonymous):

Yunika downloaded four applications for $2.99 each to her cell phone. The remaining applications that Yunika purchased cost $1.99 each. If she downloaded x total applications, which of the following equations can be used to find how much Yunika spent on cell phone application downloads?

OpenStudy (anonymous):

y = 11.96(x - 4) + 1.99 y = 11.96x + 1.99 y = 1.99(x - 4) + 11.96 y = 1.99x + 11.96